Circular No. 7077

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998fb IN ANONYMOUS GALAXY D. Reiss reports the discovery by the Mount Stromlo Abell Clus-

ter Supernova Search Team (cf. IAUC 6639) of an apparent supernova (V = 18.4) on V and R CCD images taken on Dec. 20 (limiting mag V = 19.5) by R. Hola at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998fb is located near Abell 3128 at α = 3^{h}30^{m}45^{s}.86, δ = −52^{o}32′12′′.2 (equinox 2000.0), which is 0′′.3 east and 0′′.1 north of the host galaxy’s center. A nearby star has position end figures 50^{s}.46, 31′31′′.2. SN 1998fb was not visible on similar images taken on Nov. 29 (limiting mag V = 20), but it was confirmed by Hola with the same telescope on Dec. 23 (limiting mag V = 19.1). Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

SUPERNOVA 1998fa in UGC 3513 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1998fa was taken by P. Berlind on Dec. 27.2 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), showing it to be a type-II supernova near maximum light. The spectrum exhibits a blue continuum with a broad P-Cyg Hα profile and a strong Hβ absorption. The recession velocity of the host galaxy is 7450 km/s, measured via narrow Hα emission, and the supernova photospheric expansion velocity is 13 300 km/s, measured with the broad Hα trough.

SUPERNOVA 1998es IN NGC 632 Magnitude estimates (cf. IAUC 7070): Dec. 12.510 UT, 14.6 (K. Kado-

ta and S. Yoshida, Ageo City, Saitama, Japan; unfiltered CCD); 20.822, 14.2 (M. Reszelski, Szamotuly, Poland; visual).

COMET C/1998 M5 (LINEAR) Total visual magnitude estimates: Nov. 24.72 UT, 10.0 (M. Lehký,

Hradec Králové, Czech Republic, 25×100 binoculars); Dec. 6.80, 10.0 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 11.99, 9.8 (J. Bortle, Stormville, NY, 0.41-m reflector); 15.71, 9.7 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 20.75, 9.7 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 28.71, 9.8 (Lehký; moonlight).

1998 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green